# Decentralized Coordination of Distributed Energy Resources through Local Energy Markets and Deep Reinforcement Learning

### v1

@@ 5 Conclusion
The rapid proliferation of DERs at the grid edge has led to a significant increase in variability and variance in community net load, posing challenges to electricity grid operability. In response, there has been a growing interest in TE-based DR, facilitated by community LEMs, as a viable solution to align the interests of electricity end-users and grid stakeholders [8, 31, 36]. At the same time, insights from DR system pilots highlight the necessity for automation to ensure robust participation across DR initiatives [5, 4]. In response to the decentralized and distributed nature of this challenge, model-free control approaches, particularly DRL, have emerged as promising candidates [14], fueling the interest in studies investigating the automation of participation in LEMs via DRL methods [19, 22, 23, 21, 18, 17]. While prior research has predominantly focused on socioeconomic metrics and community net load consumption, there remains a gap in demonstrating a clear reduction in variability or variance.
@@
This article addresses the research gap by utilizing a shared experience [42], recurrent PPO [44] algorithm with several modifications [48, 46, 47] to train a set of DRL agents within the context of ALEX, an economy-driven LEM where participants aim to selfishly minimize bills without information sharing [35].The trained DRL agents are compared against benchmark approaches, including a building-level DR strategy and a near-optimal dynamic programming-based solution [38]. Performance is evaluated using a set of metrics capturing net load variance across multiple time horizons, encompassing ramping rate, daily and monthly load factor, peak and average daily import and export. The experiments reveal a clear correlation between relative bill reduction and improvements in the investigated metrics. The trained DRL agents demonstrate promising performance, nearing and, in some instances, surpassing the benchmarks set by the near-optimal approach, while consistently outperforming the building-level DR strategy.
Future research directions should focus on designing more sophisticated DRL algorithms explicitly tailored to the mixed-form stochastic game nature of LEMs like ALEX. The goal is to establish a clearer performance ceiling for such solutions. Additionally, extending this investigation to diverse LEM designs could offer insights into the factors influencing the efficacy of incentivizing desired behaviors within these systems [24, 25].
Appendix A Hyperparameters
@@
The algorithm adaptations, design, and hyperparameter choices underwent testing across increasingly complex versions of the experiments discussed in the main body of this article until the performance detailed in the discussion section was achieved. The testing progression began with artificial load profiles, aiming to optimize net billing, then advanced to optimizing net billing on the City Learn dataset for a singular month, then the full year, and finally transitioned to the target application.
The advantage of this iterative process lies in the clearly defined optimal returns for the test scenarios. Recurrent PPO variants seem to vary across implementations, as the exact nature of making PPO recurrent is up to interpretation. We refer to Pleines et al. [51] for an investigation into the characteristics and sensitivities of recurrent PPO. Tests commenced with the default SB3 recurrent PPO in a shared experience replay setting [42], followed by the implementation of R2D2 [46], then state recalculation [47], and finally incorporating a learning rate schedule [48]. Each implementation underwent testing over a small range of hyperparameters for three runs each to ensure consistency, leading to the crystallization of the hyperparameter set used in this study.
@@ Acknowledgment
This research has been supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada grant RGPIN-2017-05866 and by the NSERC/Alberta Innovates grant ALLRP 561116-20.

### v2

@@ 5 Conclusion
The rapid proliferation of DERs at the grid edge increases variability and variance in community net load and challenges electricity grid operations. TE-based DR, facilitated by community LEMs, is emerging as as a viable solution by aligning the interests of electricity end-users and grid stakeholders [8, 33, 37]. At the same time, insights from DR system pilots highlight the importance of automation [5, 4]. In response to the decentralized and distributed nature of this challenge, model-free control approaches, particularly DRL, have emerged as promising candidates for the automation of participation in LEMs [19, 22, 23, 21, 18, 17]. While such prior research predominantly focused on socioeconomic metrics and community net load consumption, there remains a gap in demonstrating a clear reduction in variability or variance.
@@
This article addresses the research gap by utilizing a shared experience [43], recurrent PPO [45] algorithm with several modifications [49, 47, 48] to train a set of DRL agents within the context of ALEX, an economy-driven LEM where participants aim to selfishly minimize bills without information sharing [28].The trained DRL agents are compared against benchmark approaches, including a building-level DR strategy and a near-optimal dynamic programming-based solution [29]. Performance is evaluated using a set of metrics capturing net load variance across multiple time horizons, encompassing ramping rate, daily and monthly load factor, peak and average daily import and export. The experiments reveal a clear correlation between relative bill reduction and improvements in the investigated metrics. The trained DRL agents demonstrate promising performance, nearing and, in some instances, surpassing the benchmarks set by the near-optimal approach based on dynamic programming, while consistently outperforming the building-level DR strategy.
Future research directions should focus on exploring the transferability, generalization, and scalability of the presented framework. This includes establishing a clearer performance ceiling by designing more sophisticated DRL algorithms explicitly tailored to the mixed-form stochastic game nature of LEMs like ALEX. In this contribution relies on RL agents to learn to maximize reward. However, it would be possible to use the Constrained MDP formulation [51] to help ensure constraints are not violated. Other approaches include safety-aware RL [52] and using a controller to intervene (e.g., Learning from Intervention [53]) if the agent selects a dangerous action. Additionally, investigating the framework’s scalability should address challenges such as increased computational demands and coordination across larger agent networks, as well as exploring interactions between multiple ALEX systems under varying grid conditions. Along these lines, ALEX’s capability to integrate additional types of DERs, such as electric vehicles and renewable energy sources like wind, should also be explored, potentially leading to broader environmental benefits such as reduced emissions. The transfer of the resulting technology platform will also need to consider the diverse landscape of existing grid stakeholder regulations, alongside potential social impacts.
Additionally, extending this study’s benchmarking effort to diverse LEM designs and different agent heuristics could offer insights into the factors influencing the efficacy of incentivizing desired behaviors within these systems [24, 25].
Appendix A Auxiliary Information regarding Algorithm Setup
@@
The algorithm adaptations, design, and hyperparameter choices underwent testing across increasingly complex versions of the experiments discussed in the main body of this article until the performance detailed in the discussion section was achieved. The testing progression began with artificial load profiles, aiming to optimize net billing, then advanced to optimizing net billing on the City Learn dataset for a singular month, then the full year, and finally transitioned to the target application.The advantage of this iterative process lies in the clearly defined optimal returns for the test scenarios.
The observation space is the result of a deliberate addition of features that notably improved agent performance across these tests.
PPO algorithms across implementations vary, as discussed by Huang et al [54]. This can be extended to PPO, as the exact nature of making PPO recurrent is up to interpretation. We refer to Pleines et al. [55] for an investigation into the characteristics and sensitivities of recurrent PPO. Tests commenced with the default SB3 recurrent PPO in a shared experience replay setting [43], followed by the implementation of R2D2 [47], then state recalculation [48], and finally incorporating a learning rate schedule [49]. Each implementation underwent testing over a small range of hyperparameters for three runs each to ensure consistency, leading to the crystallization of the hyperparameter set used in this study.
@@
The training of the DRL agents was conducted on a machine with the following specifications:
  • 1.
    CPU: Ryzen 5950X
  • 2.
    GPU: Nvidia 3090
  • 3.
    RAM: 64 GB
  • 4.
    Storage: 2TB NVMe SSD
The codebase was optimized for Python 11, with the dataset stored in a SQL database for future scalability. A single run of parallel workers learning for 120 episodes took approximately 7 hours. However, it is important to note that runtimes may vary significantly depending on hardware configurations, software environments, and other factors.
@@ Acknowledgment
This research has been supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada grant RGPIN-2017-05866 and by the NSERC/Alberta Innovates grant ALLRP 561116-20. Part of this work has taken place in the Intelligent Robot Learning (IRL) Lab at the University of Alberta, which is supported in part by research grants from the Alberta Machine Intelligence Institute (Amii); a Canada CIFAR AI Chair, Amii; Digital Research Alliance of Canada; Huawei; Mitacs; and NSERC.